OpenStudy (anonymous):

Part 1 Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ 1. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. 2. Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. 3. Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C.

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (yanasidlinskiy):

Launch area (1,2) point A (1,5) first we will find the slope slope(m) = (y2 - y1) / (x2 - x1) (1,2) x1 = 1 and x2 = 2 (1,5) x2 = 1 and y2 = 5 now lets sub m = (5 - 2) / (1 - 1) m = 3/0 = 0 This is a vertical line and has an undefined slope Therefore, if we use (1,2) the equation is x = 1 if we use (1,5), the equation is x = 5 standard form : x + 0y = 5 ================================ Point A = (1,5) Point B = (2,4) find the slope m = (y2 - y1) / (x2 - x1) m = (4 - 5) / (2 - 1) m = -1/1 which is the same as -1 now put it in y - y1 = m(x - x1 form using -1 as m and either of your points. I will use (1,5) y - 5 = -1(x - 1) <-- point slope form ============================== your looking for slope intercept form y = mx + b First find the slope using points (2,4) and (4,3) Once you have found the slope(m), sub it and either points into this formula : y - y1 = m(x - x1) and just solve for x putting it into slope intercept form. ================================ I used (1,5) in question 2, so we need to use (2,4) plug in m = -1 and (2,4) x1 = 2 and y1 = 4 into this equation : y - y1 = m(x - x1) you do not need to solve it, just sub in your values ================================= now we need to solve the equations we found in problems 2 and 4 y - 5 = -1(x - 1) -- solve and put in y = mx + b form y - 4 = -1(x - 2) -- solve and put in y = mx + b form ================================ Either point you choose to use, when solved and put in y = mx + b form, will be the same answer. It does not matter what point you pick. ================================ 7 and 8.....sorry ...don't know them :)

OpenStudy (anonymous):

lol i seen that post before if u actually look you can see there 2 diffrent problems lol